Suppose the demand for money is L=0.20Y, the money supply is 200, consumption:
C=90+0.80YD, taxes T=50, Investment: I=140-5r, and government purchases: G=50.
i) Derive the IS and LM equations
1
Expert's answer
2020-08-18T12:45:26-0400

Solution:

IS equation is derived from :

"Y=C+I+G"

"Y=90+0.80YD+140-5r+50"

"Y=90+0.80(Y-T)+140-5r+50"

"Y=90+0.80Y-0.80(50)+140-5r+50"

"Y=90+0.80Y-40+140-5r+50"

"Y=0.80Y+240-5r"

"Y-0.80Y=240-5r"

"0.02Y=240-5r"

"Y=1200-25r"

"IS \t \\; equation: Y=1200-25r"


LM equation:

equate Money supply with Money demand

"(M\/P)^{s} = (M\/P)^{d}"

"200=0.20Y"

"0.20Y=200"

"Y=\frac{200}{0.20}"


"Y= 1,000"

"LM \t \\;equation: Y=1,000"